OpenStudy (anonymous):

find the exact value of the remain functions please help Cos 4/5, 270

OpenStudy (jmark):

Given Cos Ө = 4/5, 270<angle<360 => Cos^2 Ө = 4^2/5^2 => Cos^2 Ө = 16/25 we have =>sin^2 Ө +cos^2 Ө = 1 =>sin^2 Ө =1-cos^2 Ө =>sin^2 Ө=1-16/25 =>sin^2 Ө = 9/25 =>sin Ө =3/5 =>tan Ө =sin Ө /cos Ө =>tan Ө =[3/5]/[4/5] =>tan Ө =3/4 CBSE board exam date sheet http://goo.gl/Uvv2lz